OpenStudy (anonymous):

Give the next two numbers in the pattern. 4, 8, 12, 16,...

OpenStudy (anonymous):

A. 18, 20 B. 20, 24 C. 32, 64

OpenStudy (amistre64):

a good trick to see a pattern is to subtract or divide by the first number to see if it simplifies

OpenStudy (anonymous):

Add on 4 4+4=8 8+4=12 12+4=16 16+4=20 20+4=24

OpenStudy (amistre64):

if we subtract 4 from the given set: {4, 8, 12, 16} - 4 = {0, 4, 8, 12} which suggests its adding 4 each time ------------------------------ if we divide 4 from the given set: {4, 8, 12, 16} /4 = {1, 2, 3, 4} gives a good idea that these are the 4 times tables